OpenStudy (anonymous):

how do you solve a equation like this one 7^3 * 7^-6

OpenStudy (anonymous):

multiply both

OpenStudy (anonymous):

if we multiply to terms having the common base in that case we have to add their powers

OpenStudy (anonymous1):

You can use the following property: \[x^a*x^b=x^{a+b}\]

OpenStudy (anonymous):

7^(3-6)

OpenStudy (anonymous):

that is not an equation and therefore you do not solve it. what you have is an expression. this you can simplify.

OpenStudy (anonymous):

Okay so how you simplify it ?

OpenStudy (anonymous1):

\[7^3*7^{-6}=7^{3-6}=7^{-3}=\frac{1}{7^3}\]

OpenStudy (anonymous):

exactly like that...

OpenStudy (anonymous):

Thankss :)

OpenStudy (anonymous1):

You're welcome.